OpenStudy (yanasidlinskiy):

I need serious help on this! What is the inverse of the given function? y = 7x^2 - 3.

OpenStudy (mimi_x3):

isnt inverse when you switch the x and y

OpenStudy (astrophysics):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (mimi_x3):

x = 7y^2 - 3

OpenStudy (mimi_x3):

\[7y^2 = x + 3 \] \[y^2 = \frac{x+3}{7} => \]

OpenStudy (anonymous):

\[\sqrt{(1/7)y+(3/7)},-\sqrt{(1/7)y+(3/7)}\]

OpenStudy (yanasidlinskiy):

I was thinking like wouldn't it be: \[x=\pm \sqrt{\frac{ x+3 }{ 7 }}\]

OpenStudy (mimi_x3):

bingo! but should be y =....

OpenStudy (yanasidlinskiy):

So i'm correct. Right?

OpenStudy (mimi_x3):

ya

OpenStudy (yanasidlinskiy):

Thanx guys!!:D I appreciate your help:D I hate doing these. They seriously give me a hard time.